------------------------------------------------------------ Date: Wed Aug 8 08:04:56 2001, Shift: 00:00-08:00 Supervisor: JIJ, with: AJ Log book#7 pages 46 - 47 Shift Summary: ------------------------------------------------------------ 1. Beam stored and cogged since 2:15 2. Runs 4055, 4064 and 4066 with a total of ~450k events, 4066 still running. Ramiro suggested putting T1 at 1150V. This was done when beam was cogged. 3. T1 anodes tripped several times, but was set to 1150V every time. Later gating grid A on T1 trigged and we were not able to get it up to 180V again. May therefore have a broken gating grid wire. Problems with fb3 and IoServer communication. Resolved by restarting IoServer from DAQ control. ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Wed Aug 08 2001 - 08:06:01 EDT